OpenStudy (anonymous):

Two players, Hazel and Ross, play a game with a biased dice and a fair dice. Hazel chooses one of the two dice at random and rolls it. If the score is 5 or 6 she wins a point. Hazel chooses a dice, rolls it and wins a point. Find the probability that she chose the biased dice. 1=1/18 2=1/9 3=1/9 4=1/9 5=1/9 6=1/2

OpenStudy (anonymous):

See my response from earlier for help http://openstudy.com/users/order#/updates/4ee31fade4b0a50f5c5669af

OpenStudy (anonymous):

I saw that one... Am not sure how to use it to answer this question though.

OpenStudy (aravindg):

dalvoron can u hlp me after this?

OpenStudy (anonymous):

Hrm, it is a bit tougher alright. I think what you need to do is compare the P(5,6|biased) and P(5,6|unbiased). Their sum have to be 1 in this case, since we already know she won the round (i.e. rolled a 5 or a 6).

OpenStudy (anonymous):

Ok, I'll try that out and tell you if I get the answer.

OpenStudy (aravindg):

dalvoron??

OpenStudy (anonymous):

No, never mind, that's wrong.

OpenStudy (anonymous):

I haven't gotten the answer... Yeh I thought so.

OpenStudy (anonymous):

I'm completely lost, sorry!

OpenStudy (anonymous):

Hmmm... Ok. Do you know anyone else who may be able to help?

OpenStudy (anonymous):

It's hard to explain my logic here, but look at the ratio of getting a 5/6 on the two dice: \[biased : unbiased = \frac{1}{9}+\frac{1}{2}:\frac{1}{6}+\frac{1}{6}\]\[=\frac{11}{18}:\frac{6}{18}\]\[=11:6\]So the odds of her having picked the biased die is 11:6, so the probability is \[\frac{11}{11+6}=\frac{11}{17}\]

OpenStudy (anonymous):

I understand every part up to, Why do you add 11 to 6 in the denominator, if the ration is 11/6?

OpenStudy (anonymous):

11:6 means that in 11 scenarios, A happens, and in 6 scenarios, B happens. How many scenarios are there in total? Well there's 11+6=17. What is the probability that A happens? Well it's 11/17.

OpenStudy (anonymous):

It probably makes more sense in two steps like that.

OpenStudy (anonymous):

It does make a bit more sense, but I don't think I fully grasped it. I understand the ratio concept, but I still don't understand the last step...

OpenStudy (zarkon):

Here is a more traditional way to do it... \[P(B|W)=\frac{P(B,W)}{P(W)}=\frac{P(W|B)P(B)}{P(W)}\] \[=\frac{P(W|B)P(B)}{P((W\cap B)\cup (W\cap B'))}\] \[=\frac{P(W|B)P(B)}{P(W\cap B)+P(W\cap B')}\] \[=\frac{P(W|B)P(B)}{P(W|B)P(B)+P(W|B')P(B')}\] \[=\frac{\frac{11}{18}\times\frac{1}{2}}{\frac{11}{18}\times\frac{1}{2}+\frac{2}{6}\times\frac{1}{2}}=\frac{11}{17}\]

OpenStudy (anonymous):

Thanks, I understand it now :)